OpenStudy (anonymous):

the heights of a certain group of adult parrots was found to be normally distributed. The mean height is 35 cm with a standard deviation of 7cm. In a group of 1000 of these birds, how many would be more than 28 cm tall?

Miracrown (miracrown):

The first thing we need to do is to find the probability of bird being more than 28 cm tall P(X > 28) where X is the height of a parrot so we are given that the heights are normally distributed so we can use the standard normal distribution to solve this but in order to do so, we must first get the number 28 into "standard" form, that is we need to find the Z score of 28

Miracrown (miracrown):

that is done by subtracting from 28 the mean and divided by the standard deviation

Miracrown (miracrown):

 so what I did was take 28 and from it subtract the mean which is 35 and then divided by the standard deviation which is 7 so 28 - 25 = - 7 and -7/7 = -1 so we end up with P(Z > - 1) now we can get rid of the negative sign by switching > to < so P(Z > -1) = P(Z < 1) so in order to find P(Z < 1), we need to make use of a standard normal table (a Z table) do you have one? @c3darkstar